# Improving probabilistic forecasts of extreme wind speeds by training statistical post-processing models with weighted scoring rules

### v2

@@ 1 Introduction
Extreme weather events such as heavy precipitation or windstorms have the potential to cause tremendous damage to lives and livelihoods. To mitigate their impact the development of early warning systems is crucial, which rely on accurate forecasts of relevant extremes. In this work we aim to improve statistical and machine learning models for probabilistic predictions of extreme wind speeds for applications to the post-processing of numerical weather prediction (NWP) wind speed forecasts.
@@
Weather forecasts are usually generated by ensembles of NWP models, each with different initial conditions to quantify the uncertainty present in atmospheric phenomena. These ensembles, however, often contain biases and errors in dispersion, thus they necessitate the application of statistical post-processing techniques to generate accurate and well-calibrated probabilistic forecasts. This can be especially relevant for extreme events, whose forecasts might be improved most from post-processing, but which also require special care when using common techniques (Friederichs and Thorarinsdottir, 2012; Williams et al., 2014; Friederichs et al., 2018; Hess, 2020; Velthoen et al., 2023).
@@
For post-processing of wind speed forecasts with particular focus on extremes, different models have been proposed in the literature. Lerch and Thorarinsdottir (2013) suggest post-processing based on ensemble model output statistics (EMOS, Gneiting et al., 2005) – a commonly applied post-processing technique – using the truncated normal and generalized extreme value (GEV) distributions as well as a regime switching model. Baran and Lerch (2015) propose post-processing based on the log-normal and truncated normal distributions as well as regime switching between the two. Allen et al. (2021a) incorporate weather regimes based on the North Atlantic Oscillation into their wind speed post-processing based on truncated logistic distributions, showing that this leads to improved predictive performance at extremes. Oesting et al. (2017) introduce post-processing of wind gusts forecasts based on a bivariate Brown-Resnick process.
@@
Whilst the above authors propose different model structures that enable more accurate inference, we follow a different approach and change the training procedure used to fit models, in order to improve performance at extremes. This can in future work be readily extended to different model specifications. In order to improve the performance of predictive models at extremes, in recent times, a number of different authors have suggested to adjust loss functions used for the training of these models, which are often based on deep neural networks (Ding et al., 2019; Chen et al., 2022; Hess and Boers, 2022; Scheepens et al., 2023; Olivetti and Messori, 2024). However these works focus mostly on deterministic forecasts and as we show in section 3 the weighting schemes that these authors propose might not be suitable for probabilistic forecasts as they lead to "hedging" of the forecast distribution.
@@
We propose adjusting training by estimating models using weighted proper scoring rules, in particular the so-called threshold-weighted continuous ranked probability score (twCRPS). This proper scoring rule, introduced by Gneiting and Ranjan (2011) has long been used for evaluation of predictive models, but to the authors’ knowledge not yet for systematic parameter optimization. In this work we focus on post-processing using Ensemble Model Output Statistics (EMOS), a typical post-processing method, and introduce the usage of the twCRPS to estimate the model parameters. We show that this improves predictive performance for extremes, for a variety of extremal thresholds. We set out some of the choices and challenges that forecasters need to consider when doing this, and describe the effects of these choices in real and simulated data, closing with recommendations of what to consider when applying this in operational settings. We also give different characterizations of the twCRPS and derive closed-form expressions for a number of predictive distributions.
@@
This article is organized as follows. First, in section 2 we will present the data and methods used for post-processing in this study. We will then consider a motivating example in section 3 to highlight the need for propriety of scores used for the training of probabilistic forecasting models for extremes. In section 4 we introduce weighted scoring rules and the threshold-weighted continuous ranked probability score (twCRPS) that we use in section 5.1 to train post-processing models with a focus on extreme wind speeds. We show that this improves forecast performance for extremes and discuss some choices and challenges that forecasters might encounter in sections 5.1, 5.2 and 5.3, namely the body-tail trade-off, the role of the threshold for training models as well as sample based estimation. In section 6 we propose two strategies to balance different objectives that forecasters might encounter when wanting to improve forecasts for extremes: how to balance the forecast performance for the distribution body and the distribution tails. Finally, in section 7 we present a number of synthetic experiments to better understand the training impact of the twCRPS and interpret our results. We conclude in section 8, discussing the results and giving some recommendations for forecasters wishing to apply the methodology. In the appendix we give a number of different characterizations of the twCRPS and derive closed-form expressions for a number of distributions.
@@ 2 Data and methods
In this section we briefly describe the data and methods used for post-processing of wind speed forecasts.
@@ 2.1 Data
In this study we consider forecasts of 10m wind speed at surface observation stations in the United Kingdom at a lead time of 48h, initialized between 1 April 2019 and 31 March 2022. The forecasts are taken from the $0$ UTC initialization of the Met Office global ensemble prediction system (MOGREPS-G, Walters et al. 2017; Porson et al. 2020), an 18-member ensemble (including the unperturbed control run) with a horizontal resolution of 20km. The forecasts are bilinearly regridded to the locations of 124 surface synoptic observation (SYNOP) stations in the United Kingdom (see Figure 1 for the locations) and verified against wind speed observations (10min average of 10m wind speed) there. We assume the forecast ensemble to be exchangeable and summarize it by the ensemble mean and ensemble standard deviation. We use forecasts issued between 1 April 2019 and 31 December 2020 for training and forecasts issued between 1 January 2021 and 31 March 2022 as an independent test set.
Figure 1: Locations of 124 SYNOP stations used for the verification of wind speed forecasts.
2.2 Methods
@@
$\mu=\alpha+\beta m+\lambda_{\mu,{\rm sin}}\sin\left(\text{ndoy}\right)+\lambda_{\mu,{\rm cos}}\cos\left(\text{ndoy}\right),$ (2)
$\log\sigma=\gamma+\delta s+\lambda_{\sigma,{\rm sin}}\sin\left(\text{ndoy}\right)+\lambda_{\sigma,{\rm cos}}\cos\left(\text{ndoy}\right).$ (3)
Here $\mu$ and $\sigma$ are the location and scale parameters of the predictive distribution $\mathcal{D}$.
EMOS models are usually fit by either minimizing the negative log-likelihood or the continuous ranked probability score (CRPS). For a probabilistic forecast given through a cumulative distribution function (CDF) $F$ and an observation $y$ the latter is defined as:
@@
Both the CRPS and the negative log-likelihood are proper scoring rules (Gneiting and Raftery, 2007), meaning that in expectation they are minimal when the forecast distribution is the "true" distribution (see section 4 for some more discussion of propriety). Closed-form expressions of the CRPS exist for the truncated normal and truncated logistic distributions (see Jordan et al., 2019) and are used for the minimization. In this study EMOS models are fit separately for each site.
@@
Unless otherwise stated in this paper we use a Broyden-Fletcher-Goldfarb-Shanno (BFGS) quasi-Newton optimizer (Nocedal and Wright, 2006) with empirical gradient estimates for optimizing scoring functions, which falls back onto a Nelder-Mead optimizer (Nelder and Mead, 1965) should it fail. We follow the practice of the crch package (Messner et al., 2022) and use the coefficients of a linear regression as starting values for the location parameters ($\alpha,\beta,\lambda_{\mu,{\rm sin}},\lambda_{\mu,{\rm cos}}$) and the exponentially transformed standard error as the starting value of the intercept of the scale parameter ($\gamma$), whilst setting all other scale parameters ($\delta,\lambda_{\sigma,{\rm sin}},\lambda_{\sigma,{\rm cos}}$) initially to zero.
3 Motivating the use of proper scoring rules
@@
In order to improve the performance of data-driven models for prediction of extreme events multiple authors propose adjusting loss functions used to learn parameters of these models (e.g. Ding et al., 2019; Chen et al., 2022; Olivetti and Messori, 2024). Among others Hess and Boers (2022) and Scheepens et al. (2023) propose weighting schemes for loss functions to weight misprediction at extremes stronger during training. Many of the above authors focus on deterministic predictions of extremes, however as Friederichs and Thorarinsdottir (2012) point out to allow for the assessment of the potentially large uncertainty in predictions of extreme events, such predictions should be "probabilistic in nature". In this section we will study the effect of a loss function weighting strategy for probabilistic extreme event forecasting in the setting of post-processing. This is to motivate the need for proper scoring rules when training probabilistic forecasting models.
@@
Hess and Boers (2022) propose a weighting scheme for loss functions to penalize strongly the mis-prediction of high precipitation values during training. They only consider deterministic forecasts, trained using root mean squared error (RMSE), however we consider this weighting scheme adapted for the training of probabilistic post-processing models for wind speed. We consider EMOS post-processing models using a truncated normal distribution, trained by minimizing the following loss (henceforth referred to as HB-loss):
@@
This weighting scheme means that prediction quality (measured through the CRPS) of high observed wind speed values will carry more weight during training. We choose $a$ as the location-wise 80th percentile of the observations of the training dataset and $b$ as the location-wise standard deviation of the observational training dataset. We score predictions against EMOS models trained classically using the CRPS.
Investigating the trained model’s performance on the test set we find that when training EMOS post-processing models using the HB-loss the resulting models have much lower HB-loss on the test set (16% lower) compared to the CRPS-trained model. Looking at the dataset of test set extremes – wind speed values over the location-wise 80th percentile of the training set $\tau$ – we find that the HB-trained model is far better than the CRPS-trained model having 26% better RMSE and 6% better CRPS respectively on this dataset. The model has good out-of-sample HB-loss performance and seems to be able to capture the intensity of extremes quite well.
@@
$BS_{\tau}(F,y):=\left[F(\tau)-\mathbb{1}\{y\geq\tau\}\right]^{2}.$ (6)
We find that the HB-trained model strongly overpredicts extremes, having 12% worse Brier Score than the CRPS-trained model. A similar picture can be seen by considering probability integral transform (PIT) histograms (Thorarinsdottir and Schuhen, 2018). These are histograms of the predictive CDF evaluated at the observation and should appear uniform under calibration. Figure 2 shows the histograms for the CRPS and HB-trained model. Whilst the CRPS-trained model seems approximately calibrated the HB-trained model has a right tail that is much too heavy and seems to strongly underpredict values in the body of the climatological distribution.
Figure 2: PIT histograms for wind speed post-processing models using truncated normal distributions, trained using CRPS loss (left) and HB-loss (eq. 5, right).
When evaluating only on test set extremes the HB model seems to be a suitable prediction model, however when considering the full dataset the model strongly overpredicts the occurrence of extremes and seems strongly mis-calibrated. A similar result also occurs with a very related strategy of oversampling extremes prior to training, often used to adjust for class imbalance in classification problems (van den Goorbergh et al., 2022). The need to evaluate not only the intensity of extreme event forecasts, but to do this jointly with the predicted occurrence has been highlighted by many authors and is commonly referred to as the "forecaster’s dilemma" (Lerch et al., 2017). If only forecasts of intensity are evaluated, then this encourages forecasters to overpredict extreme values, something that is not a problem in cases where there is no penalty for false-positive extreme event warnings. However in real applications this will quickly erode trust in any forecasting system and lead to economic costs. The CRPS is a proper scoring rule, thus using it for training and evaluation encourages honesty of the forecasters and calibrated forecasts. However when weighting the CRPS by an observed outcome, for example by restricting evaluation to the dataset of test set extremes or by using a weighting such as the HB-weighting, propriety is lost. This makes the score unsuitable for objective evaluation and for training (see next section).

### v4

@@ 1 Introduction
Extreme weather events such as heavy precipitation or windstorms have the potential to cause tremendous damage to lives and livelihoods. To mitigate their impact the development of early warning systems is crucial, which rely on accurate forecasts of relevant extremes. In this work, we aim to improve statistical and machine learning models for probabilistic predictions of extreme wind speeds for applications to the post-processing of numerical weather prediction (NWP) wind speed forecasts.
@@
Weather forecasts are usually generated by ensembles of NWP models, each with different initial conditions and possibly perturbed models, to quantify the uncertainty present in atmospheric phenomena. These ensembles, however, often contain biases and errors in dispersion, thus necessitating the application of statistical post-processing techniques to generate accurate and well-calibrated probabilistic forecasts. This can be especially relevant for extreme events, whose forecasts might be improved most from post-processing, but which also require special care when using common techniques (Friederichs and Thorarinsdottir, 2012; Williams et al., 2014; Friederichs et al., 2018; Hess, 2020; Velthoen et al., 2023).
@@
For post-processing of wind speed forecasts with particular focus on extremes, different models have been proposed in the literature. Lerch and Thorarinsdottir (2013) suggest post-processing based on ensemble model output statistics (EMOS, Gneiting et al., 2005) – a commonly applied post-processing technique – using the truncated normal and generalized extreme value (GEV) distributions as well as a regime switching model. Baran and Lerch (2015) propose post-processing based on log-normal and truncated normal distributions as well as regime switching between the two. Allen et al. (2021a) incorporate weather regimes based on the North Atlantic Oscillation into their wind speed post-processing using truncated logistic distributions, showing that this leads to improved predictive performance at extremes. Oesting et al. (2017) introduce post-processing of wind gusts forecasts based on a bivariate Brown-Resnick process.
@@
Whilst the above authors propose different model structures that enable more accurate inference, we follow a different approach and change the training procedure used to fit models, in order to improve performance at extremes. This can in future work be readily extended to different model specifications. In order to improve the performance of predictive models at extremes, in recent times, a number of different authors have suggested to adjust loss functions used for the training of these models, which are often based on deep neural networks (Ding et al., 2019; Chen et al., 2022; Hess and Boers, 2022; Scheepens et al., 2023; Olivetti and Messori, 2024). However, these works focus mostly on deterministic forecasts, and as we show in section 3 the weighting schemes that these authors propose might not be suitable for probabilistic forecasts as they lead to hedging of the forecast distribution.
@@
We propose adjusting training by estimating models using weighted proper scoring rules, in particular the so-called threshold-weighted continuous ranked probability score (twCRPS). This proper scoring rule, introduced by Gneiting and Ranjan (2011) has long been used for evaluation of predictive models, but to the authors’ knowledge not yet for systematic parameter optimization. In this work, we focus on post-processing using ensemble model output statistics (EMOS), a typical post-processing method, and introduce the usage of the twCRPS to estimate the model parameters. We show that this improves predictive performance for extremes, for a variety of extremal thresholds. We set out some of the choices and challenges that forecasters need to consider when doing this, and describe the effects of these choices in real and simulated data, closing with recommendations of what to consider when applying this in operational settings. We also give different characterizations of the twCRPS and derive closed-form expressions for a number of predictive distributions.
@@
This article is organized as follows. First, in section 2 we will present the data and methods used for post-processing and forecast verification in this study. We will then consider a motivating example in section 3 to highlight the need for propriety of scores used for the training of probabilistic forecasting models for extremes. In section 4 we introduce weighted scoring rules and the threshold-weighted continuous ranked probability score (twCRPS) that we use in section 5 to train post-processing models with a focus on extreme wind speeds. We show that this improves forecast performance for extremes and discuss some choices and challenges that forecasters might encounter in sections 5.1, 5.2 and 5.3, namely the role of the threshold for training models and the role of the training datasize. In section 6 we propose two strategies to balance different objectives that forecasters might encounter when wanting to improve forecasts for extremes: how to balance the forecast performance for the distribution body and the distribution tails. Finally, in section 7 we present a number of synthetic experiments to better understand the training impact of the twCRPS and interpret our results. We conclude in section 8, discussing the results and giving some recommendations for forecasters wishing to apply the methodology. In the Appendix we give a number of different characterizations of the twCRPS and derive closed-form expressions for a number of distributions.
@@ 2 Data and methods
In this section we briefly describe the data and methods used for post-processing of wind speed forecasts and verification.
@@ 2.1 Data
In this study we consider forecasts of 10m wind speed at surface observation stations in the United Kingdom at a lead time of 48h, initialized between 1 April 2019 and 31 March 2022. The forecasts are taken from the $0$ UTC initialization of the Met Office global ensemble prediction system (MOGREPS-G, Walters et al. 2017; Porson et al. 2020), an 18-member ensemble (including the unperturbed control run) with a horizontal resolution of 20km. The forecasts are bilinearly regridded to the locations of 124 surface synoptic observation (SYNOP) stations in the United Kingdom (see Figure 1 for the locations) and verified against wind speed observations (10min average of 10m wind speed) there. We summarize the forecast ensemble by the ensemble mean and ensemble standard deviation. We use forecasts issued between 1 April 2019 and 31 December 2020 for training and forecasts issued between 1 January 2021 and 31 March 2022 as an independent test set.
Figure 1: Locations of 124 SYNOP stations used for the verification of wind speed forecasts, together with associated clusters (see section 2.2 and table 1 in the appendix for explanation).
2.2 Post-processing methods
@@
$\mu=\alpha+\beta m+\lambda_{\mu,{\rm s}}\sin\left(\text{ndoy}\right)+\lambda_{\mu,{\rm c}}\cos\left(\text{ndoy}\right),$ (2)
$\log\sigma=\gamma+\delta s+\lambda_{\sigma,{\rm s}}\sin\left(\text{ndoy}\right)+\lambda_{\sigma,{\rm c}}\cos\left(\text{ndoy}\right).$ (3)
Here $\mu$ and $\sigma$ are the location and scale parameters of the predictive distribution $\mathcal{D}$. Sometimes, the predictive standard deviation $s$ is also log-transformed to ensure that $s$ and $\sigma$ are on the same scale, however, this did not make a difference in performance here.
EMOS models are usually fit by minimizing either the negative log-likelihood or the continuous ranked probability score (CRPS). For a probabilistic forecast given through a cumulative distribution function (CDF) $F$ and an observation $y$ the latter is defined as:
@@
Both the CRPS and the negative log-likelihood are proper scoring rules (Gneiting and Raftery, 2007), meaning that in expectation they are minimal when the forecast distribution is the true distribution (see section 4 for some more discussion of propriety). Closed-form expressions of the CRPS exist for the truncated normal and truncated logistic distributions (see Jordan et al., 2019) and are used for the minimization.
EMOS models can be fit individually to a single site, or to sets of sites. For example, the configuration of the Met Office IMPROVER system described in Roberts et al. (2023) fits a single EMOS model for wind speed to post-process forecasts jointly to measurement sites, including the 124 stations presented here. As we will highlight in section 5.3 large data is needed for training using weighted scoring rules. Thus, to increase the datasize for training and evaluation of each EMOS fit we use a semilocal post-processing approach as introduced by Lerch and Baran (2016). In this approach EMOS models are fit to clusters of similar stations. We cluster based on the observed climatology of wind speed at each station, which has the advantage that quantile-based thresholds are homogeneous within each cluster.
Following Lerch and Baran (2016) we compute K = 24 features for each station, corresponding to the $1/(K+1),\dots K/(K+1)$ quantiles of the location-wise distribution of observed wind speed in the training dataset. In a second step k-means clustering (Murphy and Bach, 2012) is used to identify sets of similar stations from the features of all stations. We use the “elbow-method" to identify four distinct clusters, shown in Figure 1. The clusters broadly correspond to coastal locations (red), more exposed coastal locations (blue), more exposed inland locations (green) and less exposed inland locations (purple). In a second step EMOS models are then fit to jointly post-process wind speed forecasts for all stations within one cluster. Although four clusters are used here, the results below were consistent for other choices of cluster numbers, as long as a minimum amount of stations were present in each cluster. The 80th and 90th percentile wind speed in each cluster are shown in Table 1 in the appendix.
@@
Unless otherwise stated in this paper we use a Broyden-Fletcher-Goldfarb-Shanno (BFGS) quasi-Newton optimizer (Nocedal and Wright, 2006) with empirical gradient estimates for optimizing scoring functions, which falls back onto a Nelder-Mead optimizer (Nelder and Mead, 1965) should it fail. Using analytic gradients is also possible and might increase the optimization efficiency, but we leave such exploration for future work. We follow the practice of the crch package (Messner et al., 2022) and use the coefficients of a linear regression as starting values for the location parameters ($\alpha,\beta,\lambda_{\mu,{\rm s}},\lambda_{\mu,{\rm c}}$) and the log-transformed standard error as the starting value of the intercept of the scale parameter ($\gamma$), whilst setting all other scale parameters ($\delta,\lambda_{\sigma,{\rm s}},\lambda_{\sigma,{\rm c}}$) initially to zero.
2.3 Forecast verification
The goal of probabilistic forecasting has been put by Gneiting et al. (2005) as to “maximize sharpness subject to calibration". Sharpness is a measure of the forecast distribution only and refers to the concentration of the forecast density function, whilst calibration is a joint property between observations and forecast and corresponds to the consistency between the forecast distributions and the observations. As this paper focuses on forecasts of extreme events, this will also guide our forecast verification.
Sharpness of forecasts can be evaluated via statistics of the predictive distribution. More often however, this is evaluated jointly with calibration via the usage of proper scoring rules. In particular we will use the CRPS and threshold-weighted CRPS for forecast evaluation, which are both proper scoring rules, and which we will introduce in section 4. Calibration on the other hand, which is a “necessary condition for the optimal use and value of a forecast" (Thorarinsdottir and Schuhen, 2018), is usually also assessed individually. Most commonly calibration is evaluated using probability integral transform (PIT) histograms (see for example Thorarinsdottir and Schuhen, 2018). These are histograms of the predictive CDF $F_{i}$, evaluated at the observation $y_{i}$: $F_{i}(y_{i})$. Under probabilistic calibration PIT values should be uniform, corresponding to a flat histogram. Alternatively, they can also be visualized against theoretical quantiles from a uniform distribution in a quantile-quantile plot (QQ-plot), in which case the PIT values should lie on the diagonal axis.
As pointed out by Allen et al. (2023a); Gneiting and Resin (2023); Allen et al. (2024), probabilistic calibration as assessed by uniformity of PIT histograms does not necessarily imply probabilistic calibration for all subsets of the observations that are of interest. For example, a forecast could be probabilistically calibrated overall, without being probabilistically calibrated in the tail of the observational distribution. To assess probabilistic calibration for tail events one can evaluate conditional PIT histograms as suggested by Allen et al. (2023a); Mitchell and Weale (2023). For a forecast $F_{i}$ define the excess distribution over a threshold $\tau$ as:
$F_{i,\tau}(x)=\frac{F_{i}(x)-F_{i}(\tau)}{1-F_{i}(\tau)},\quad x\geq\tau$
The conditional PIT (CPIT) values are the values of the forecast excess distribution evaluated at observations $y_{i}\geq\tau$: $F_{i,\tau}(y_{i})$. These should again be uniform if the forecast is probabilistically calibrated over the threshold (in the tails) which can again be assesed using histograms or QQ-plots.
The conditional PIT assesses the probabilistic calibration for the distribution of threshold exceedances. Allen et al. (2024) also introduce the notion of probabilistic tail calibration, which in addition to the excess distribution also incorporates the probability of threshold exceedance. A forecast is tail calibrated if the following holds for $\tau$ approaching the upper endpoint of the distribution of observations e.g. $\tau\to\infty$ :
$\frac{\mathbb{P}(F_{\tau}(Y)\leq u,Y>\tau)}{\mathbb{E}(1-F(\tau))}=u\quad\text{for all}\quad u\in[0,1].$
This can be empirically assessed using the following ratio:
$\hat{R}_{\tau}(u)=\frac{\sum_{i\in I_{\tau}}\mathbb{1}\{z_{i}^{\tau}\leq u\}}{\sum_{i=1}^{n}(1-F_{i}(\tau))}=\frac{1}{|I_{\tau}|}\sum_{i\in I_{\tau}}\mathbb{1}\{z_{i}^{\tau}\leq u\}\cdot\frac{|I_{\tau}|}{\sum_{i=1}^{n}(1-F_{i}(\tau))},\quad u\in[0,1].$
here $z_{1}^{\tau},\dots z_{n}^{\tau}$ are the conditional PIT values and $I_{\tau}=\{i:y_{i}>\tau\}$. This ratio should be equal to or close to $u$ for all values of $u\in[0,1]$. It corresponds to the empirical distribution of conditional PIT values, multiplied by the occurrence ratio of realized over expected threshold exceedances. To summarize the behavior of $\hat{R}_{\tau}(u)$ the authors also suggest using the following as a measure of tail miscalibration (TMCB):
$\text{TMCB}=\sup_{u\in[0,1]}\left|\hat{R}_{\tau}(u)-u\right|.$ (5)
3 Motivating the use of weighted proper scoring rules
@@
In order to improve the performance of data-driven models for prediction of extreme events multiple authors propose adjusting loss functions used to learn parameters of these models (e.g. Ding et al., 2019; Chen et al., 2022; Olivetti and Messori, 2024). Among others Hess and Boers (2022) and Scheepens et al. (2023) propose weighting schemes for loss functions to weight misprediction at extremes stronger during training. Many of the above authors focus on deterministic predictions of extremes, however as Friederichs and Thorarinsdottir (2012) point out to allow for the assessment of the potentially large uncertainty in predictions of extreme events, such predictions should be “probabilistic in nature". In this section we will study the effect of a loss function weighting strategy for probabilistic extreme event forecasting in the setting of post-processing. This is to motivate the need for proper scoring rules, when training probabilistic forecasting models, which we discuss in the next section.
@@
Hess and Boers (2022) propose a weighting scheme for loss functions to penalize stronger the mis-prediction of high precipitation values during training. They only consider deterministic forecasts, trained using root mean squared error (RMSE), however, we consider this weighting scheme adapted for the training of probabilistic post-processing models for wind speed. We consider EMOS post-processing models using a truncated normal distribution, trained by minimizing the following loss (henceforth referred to as HB-loss):
@@
This weighting scheme means that prediction quality (measured through the CRPS) of high observed wind speed values will carry more weight during training. We choose $a$ as the cluster-wise 90th percentile of the observations of the training dataset and $b$ as the cluster-wise standard deviation of the observational training dataset. We score predictions against EMOS models trained classically using the CRPS.
Investigating the trained model’s performance on the test set we find that when training EMOS post-processing models using the HB-loss the resulting models have much lower HB-loss on the test set (9.5% lower) compared to the CRPS-trained model. Looking at the dataset of test set extremes – wind speed values over the location-wise 90th percentile of the training set $\tau$ – we find that the HB-trained model is far better than the CRPS-trained model having 86.45% better RMSE and 26.3% better CRPS respectively on this dataset. The model has good out-of-sample HB-loss performance and seems to be able to capture the intensity of extremes quite well.
@@
$BS_{\tau}(F,y):=\left[\left(1-F(\tau)\right)-\mathbb{1}\{y\geq\tau\}\right]^{2}.$ (7)
We find that the HB-trained model strongly overpredicts extremes, having 18% worse Brier score than the CRPS-trained model. A similar picture can be seen by considering PIT histograms. Figure 2 shows the histograms for the CRPS and HB-trained model. Whilst the CRPS-trained model seems approximately calibrated the HB-trained model has a right tail that is much too heavy and seems to strongly underpredict values in the body of the climatological distribution.
Figure 2: PIT histograms for wind speed post-processing models using truncated normal distributions, trained using CRPS loss (left) and HB-loss (equation 6, right). Pooled over all clusters.
When evaluating only on test set extremes the HB model seems to be a suitable prediction model, however, when considering the full dataset the model strongly overpredicts the occurrence of extremes and seems strongly mis-calibrated. A similar result also occurs with a very related strategy of oversampling extremes prior to training, often used to adjust for class imbalance in classification problems (van den Goorbergh et al., 2022). The need to evaluate not only the intensity of extreme event forecasts, but to do this jointly with the predicted occurrence has been highlighted by many authors and is commonly referred to as the “forecaster’s dilemma" (Lerch et al., 2017). If only forecasts of intensity are evaluated, then this encourages forecasters to overpredict extreme values, something that is not a problem in cases where there is no penalty for false-positive extreme event warnings. However, in real applications this will quickly erode trust in any forecasting system and lead to economic costs. The CRPS is a proper scoring rule, thus using it for training and evaluation encourages honesty of the forecasters and calibrated forecasts. However, when weighting the CRPS by an observed outcome, for example by restricting evaluation to the dataset of test set extremes or by using a weighting such as the HB-weighting, propriety is lost. This makes the score unsuitable for objective evaluation and for training (see next section).